# Unveiling the Siren’s Song: Towards Reliable Fact-Conflicting Hallucination Detection

### v1

@@
Abstract.
Large Language Models (LLMs), such as ChatGPT/GPT-4, have garnered widespread attention owing to their myriad of practical applications, yet their adoption has been constrained by issues of fact-conflicting hallucinations across web platforms. The assessment of factuality in text, produced by LLMs, remains inadequately explored, extending not only to the judgment of vanilla facts but also encompassing the evaluation of factual errors emerging in complex inferential tasks like multi-hop, and etc. In response, we introduce FactCHD, a fact-conflicting hallucination detection benchmark meticulously designed for LLMs. Functioning as a pivotal tool in evaluating factuality within “Query-Respons” contexts, our benchmark assimilates a large-scale dataset, encapsulating a broad spectrum of factuality patterns, such as vanilla, multi-hops, comparison, and set-operation patterns. A distinctive feature of our benchmark is its incorporation of fact-based chains of evidence, thereby facilitating comprehensive and conducive factual reasoning throughout the assessment process. We evaluate multiple LLMs, demonstrating the effectiveness of the benchmark and current methods fall short of faithfully detecting factual errors. Furthermore, we present Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence [^1] The benchmark dataset and source code will be made available in https://github.com/zjunlp/FactCHD..
Fact-Conflicting, Hallucination Detection, LLMs, Benchmark
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NY
1. Introduction
Figure 1. Illustrating Fact-Conflicting Hallucination Detection.
There is a significant risk that LLMs (Zhao et al., 2023) may generate web content that appears credible but is factually incorrect or imprecise (Yin et al., 2023; Ji et al., 2023; Huang et al., 2023), thereby spreading misinformation on the web. This phenomenon is known as fact-conflicting hallucination (Zhang et al., 2023) and may result from flawed or outdated knowledge incorporated into the model’s parameters, as well as the language model’s shortcomings in tasks such as fact reconciliation, logical reasoning, and other sophisticated cognitive processes. This limitation not only restricts the application of generative AI in high-stakes areas such as finance, healthcare and law but also exacerbates the dissemination of erroneous information on the web community, posing a significant obstacle to the healthy development of the web.
@@ 1. Introduction
Given their pivotal role in ensuring the security of online content, effectively detecting and mitigating fact-conflicting hallucinations stands paramount. Traditional task definitions for fact verification (Wadden et al., 2020a, b) exhibit limitations, particularly in accommodating writing tasks that users frequently navigate when interfacing with LLMs. Moreover, existing hallucination evaluation benchmarks (Li et al., 2023a; Muhlgay et al., 2023), predominantly centering on rudimentary facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal. In light of this, we introduce an inherently rigorous, yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence. For instance, in Figure 1, when confronted with a query and its generated response, evaluators, within the prescribed scenario, are impelled to harness both their intrinsic knowledge and external resources, while employing reasoning skills to render a factual judgment accompanied by an elucidative reasoning process.
In paving the way for future strides in hallucination evaluation, we dedicate our efforts towards establishing a rigorous benchmark, Fact-Conflicting Hallucination Detection (FactCHD), tailored for LLMs and encompassing a variegated array of factuality patterns, including Vanilla, Multi-hops, Comparison, and Set-Operation patterns. Drawing inspiration from the adage “to know it and to know the reason why of it” by Zhuzi, FactCHD endeavors to intertwine fact-based chains of evidence, providing cogent justifications when predicting the factual or non-factual nature of a statement. Achieving the intended goal of gathering a substantial amount of data through full-process human annotation is a time-consuming and resource-intensive task with limited scalability. Considering the high reliability of existing knowledge graphs (KGs) and textual knowledge, we advocate for a data construction strategy that leverages knowledge as facts, coupled with semi-supervised annotation, to facilitate the crafting of the aforementioned benchmark.
As illustrated in Figure 3, we engineer our benchmark through a systematic construction strategy, bifurcated into two primary sections, encompassing six meticulous steps. Initially, a subset of queries is input into several LLMs, with corresponding responses being manually annotated, thereby curating a minimal but realistic “Query-Response” dataset enriched with factuality information. Subsequently, we intertwine knowledge sub-graphs and textual knowledge, utilizing these as factual underpinnings to guide ChatGPT in the autonomous generation of contextual “Query-Response” pairs. Then human similarity assessment is employed iteratively to finetune prompts, ensuring stylistic coherence between generated and realistic factuality data. The fourth step involves automated similarity screening of the synthesized factuality data, identifying and excising highly similar samples to uphold dataset diversity. Fifthly, ChatGPT is guided by precise instructions to generate evidence chains, anchoring its judgments in the “Query-Response” contexts on solid knowledge. Concludingly, an adept team of 21 individuals meticulously filters the data, adhering to stipulated rules, thereby assuring the procurement of a high-quality dataset.
@@
We conduct baseline experiments to assess the efficacy of several LLMs (such as Alpaca, ChatGPT, etc), utilizing our FactCHD under various settings—zero-shot, in-context learning, tuning for detect-specific expertise, and knowledge enhancement through retrieval/tools. Despite tuning and knowledge augmentation positively influencing the evaluation of fact-conflicting hallucinations, both open-source LLMs and ChatGPT encounter challenges in the precise and robust detection of factual inaccuracies. Thus, we introduce Truth-Triangulator, our innovative framework, inspired by the “Triangulation” theory (Valenza, 2016). Within this structure, tool-enhanced ChatGPT is designated as the Truth Seeker, and the detect-specific expert (based on Llama2-7B-LoRA) serves as the Truth Guardian. An additional agent, the Fact Verdict Manager, is tasked with amassing evidence from various perspectives to fortify the reliability and accuracy of the derived conclusions. Truth-Triangulator accentuates employing cross-referencing generators and search capabilities to astutely evaluate and adjudicate responses with potential factual discrepancies. Key insights are summarized as follows:
  • We devise FactCHD, a comprehensive large-scale factuality evaluation benchmark that encompasses multiple domains and a myriad of factual dimensions, complemented by interpretable fact-based chains of evidence. This benchmark is poised to be an invaluable asset, propelling future innovations in detecting fact-conflicting hallucinations.
  • We introduce a semi-supervised data construction strategy, utilizing knowledge as facts to expedite the assembly of the fact-conflicting hallucination dataset. Owing to the authenticity and extensive domain coverage of KGs, this strategy not only assures enhanced scalability but also stands out as particularly adept for large-scale applications.
  • We present Truth-Triangulator, a framework employing triangulation for truth discernment, utilizing cross-referencing generators and search tools to arbitrate responses of dubious factuality. Initial experiments substantiate the varied performance of different LLMs in identifying fact-conflicting hallucinations and affirm the superiority of our proposed approach.
2. Related Work
@@
Despite the impressive capabilities of LLMs (Zhao et al., 2023; Chen et al., 2022b, a; Zhang et al., 2022), such as ChatGPT (OpenAI, 2022), in comprehending and carrying out user instructions, they sometimes generate misleading information with a high level of confidence. This phenomenon, known as hallucination (Ji et al., 2023; Ye et al., 2023), has been classified by (Zhang et al., 2023) into three categories: input-conflicting, context-conflicting, and fact-conflicting hallucinations. While humans can relatively easily identify the first two types, fact-conflicting hallucinations present a significant challenge as they involve the dissemination of misleading information on the web due to conflicting factual knowledge. While hallucination in vanilla natural language generation (NLG) settings has been extensively studied (Shuster et al., 2021; Creswell and Shanahan, 2022; Shuster et al., 2021; Das et al., 2023; Mallen et al., 2022), these methods primarily focus on analyzing hallucination across various NLP tasK.  Li et al. (2023a) recently introduced HaluEval, the hallucination evaluation benchmark evaluating the performance of LLMs in recognizing hallucinations. Different from HaLuEval which only analyzes ChatGPT’s ability to evaluate hallucination, we specifically focus on the evaluation of fact-conflicting hallucination by constructing an interpretable benchmark that can serve as a public platform for checking the factual errors of context with explanation, as well as designing the Truth-Triangulator framework for reliable hallucination detection.
@@
Table 1. Distribution of our proposed benchmark, FactCHD
Table 2. Comparison with existing fact-checking datasets. Our FactCHD include both general and vertical domains, such as health, COVID-19, climate, science, and medicine (genes, virus and disease). Van., Mul., Com. and Set. are abbreviations for the four distinct factuality patterns in FactCHD. We utilize the following abbreviations: HG - human-curated datasets; MIX - datasets containing both human-constructed and machine-generated data; Q&R - “Query-Response” context; Consist. - classification of consistency between claims and given evidence; Fact. - classification of factuality, with labels factual and non-factual; Chain. - chains of evidence.
Datasets # Source Domains Construction Format Factuality Pattern Evaluation Expandability
Fever (Thorne et al., 2018) Text General HG Claim Van. Consist.
Climate-Fever (Diggelmann et al., 2020) Text Climate Change HG Claim Van. Consist.
Health-Fever (Sarrouti et al., 2021) Text Health HG Claim Van. Consist.
Sci-Fact (Wadden et al., 2020a) Text Scientific HG Claim Van. Consist.
CoVERT (Mohr et al., 2022) Text COVID-19 HG Claim Van. Consist.
TabFact (Chen et al., 2020) Table General HG Claim Van. Consist.
HoVer (Jiang et al., 2020) Text General HG Claim Mul. Consist.
FEVEROUS (Aly et al., 2021) Text+Table General HG Claim Van. Consist.
HaluEval (Li et al., 2023a) Text General MIX Q&R Van. Hallucination
FactCHD (ours) KGs (
) & Text
General&Vertical MIX Q&R Van.&Mul.&Com.&Set. Fact.+Chain.
Factuality Detection.
Our work is also related to previous fact verification or actuality evaluation in various NLP tasks (Thorne et al., 2018; Wadden et al., 2020b; Gupta et al., 2022; Dziri et al., 2022; Rashkin et al., 2021; Kryściński et al., 2020; Pagnoni et al., 2021; Tang et al., 2022). Going beyond a simplistic binary notion of factuality, the FRANK framework (Pagnoni et al., 2021) advocates for a nuanced comprehension of factuality and introduces a typology of factual errors. Dhingra et al. (2019) measures $n$-gram lexical entailment for table-to-text generation. Recently,  Muhlgay et al. (2023) proposes FACTOR score describes LM’s tendency to assign greater likelihood to the original fact than to any of the non-factual alternatives. FELM (Chen et al., 2023) extends actuality evaluation into math and reasoning tasks. However, these benchmarks and metrics may not align well with the naturally occurring “Query-Response” format in the context of LLMs, and they emphasize accuracy evaluation while neglecting the provision of explanatory justifications for decision-making. In contrast, our work (1) provides challenge evaluation data in “Query-Response” format for LLMs; (2) considers an additional dimension of evaluation metrics that evaluate the interpretability of detecting fact-conflicting hallucination; and (3) investigates the effectiveness of Truth-Triangulator with verifying and confirming information by cross-referencing multiple independent sources or perspectives. Table 2 shows the comparison with other relevant datasets. Moreover, factuality-based hallucination detection establishes the foundation for model editing (Yao et al., 2023; Mao et al., 2023; Li et al., 2023b; Wang et al., 2023) of LLMs.
3. FactCHD Benchmark Construction
We first define the task of fact-conflicting hallucination detection and subsequently detail the benchmark construction process in the following subsections.
Task Formulation. The detection of fact-conflicting hallucinations in LLM responses in real-world scenarios entails identifying factual inaccuracies and considering potential human-initiated requests. A detector should not only predict the factuality label accurately but also generate an explanatory justification for its binary prediction. Thus, the task of fact-conflicting hallucination detection is formally defined as: Input: A question $Q$ and response $R$ from LLMs, encompassing four types of fact-conflicting: Vanilla, Multi-hop, Comparison, and Set-Operation. Output: A sequence $A=[l,e]$, where $l$ and $e$ represent the factuality label (factual or non-factual) and explanation, respectively.
Based on the aforementioned definitions, we construct FactCHD, which comprises a comprehensive array of training samples, enriched with an additional 6,960 meticulously curated samples for evaluating fact-conflicting hallucinations generated by LLMs. Our dataset ensures an equitable balance between factual or non-factual categories, providing a sturdy platform for evaluation. Notably, FactCHD presents three distinct characteristics: (1) It encompasses patterns of diverse reasoning, including multi-hop, comparison, and set operations as shown in Table 1; (2) FactCHD adheres to realistic scenarios by supplying “Query-Response” pairs and their associated evidence to validate the information provided; (3) The benchmark is meticulously crafted, leveraging KGs for initial data construction, followed by meticulous human verification to assure quality. Furthermore, the dataset inherently allows for expansion through updates based on KGs, offering a unique advantage in maintaining its contemporaneity and expandability. Further specific instructions and case data are illustrated in Appendix A. We proceed to introduce the design principles of the fact-conflicting hallucination detection benchmark in the following parts.
Figure 2. Distribution of FactCHD across various domains, with “General-p” and “General-d” denoting prevalent domains sourced from Wikipedia and Wikidata, respectively.
Addressing hallucinations in the outputs of LLMs necessitates a meticulous examination of both hallucinatory and non-hallucinatory instances. The susceptibility of LLMs to fact-conflicting hallucinations stems from their circumscribed knowledge and suboptimal reasoning capacities (Ji et al., 2023). Accordingly, we categorize four distinct patterns of factual errors, vividly illustrated in the lower section of Figure 3: (1) vanilla pattern that deals with factual statements that can be objectively verified using established sources and are typically more straightforwardly identified; (2) multi-hops pattern denotes the process of drawing conclusions by connecting multiple pieces of facts, (3) Comparison pattern refers to the act of evaluating and comparing the relative worth, and relationships between different pieces of facts, and (4) Set-operation pattern involves manipulating and combining sets of elements using operations to analyze relationships between different facts. Subsequently, we query open-source LLMs using specific instructions that reflect the identified factuality patterns. Through manual annotation of the hallucination responses, we curate a genuine compilation of hallucinations. Our objective is to gather a substantial amount of bulk hallucination data to simulate these realistic hallucinations as demonstrative examples.
3.1. Generate “Query-Response” Contexts
Capitalize knowledge as Fact
KGs, with their extensive repositories of entities and relational data, not only champion compositional reasoning but also offer a foundational structure for factual information. Concurrently, textual knowledge, while invaluable for vanilla inference, takes on a pivotal role in offering additional and nuanced information. With this in mind, our aim steers toward exploring the capabilities of utilizing existent knowledge as a factual backbone in the semi-automated assembly of fact-conflicting hallucinations, detailed as follows:
(1) We extract KG data from Wikidata (Vrandecic and Krötzsch, 2014) and PrimeKG (Chandak et al., 2023), utilizing it as an underpinning knowledge base for generating “Query-Respons” data. By manually selecting $438$ prevalent relations and crafting subgraphs through $K$-hop walks, performed $N$ times from divergent, randomly selected starting entities, we procure a varied set of subgraphs to be employed for multi-hop reasoning, fact comparison, and set operation. Heuristic rules are applied to ensure minimal intersection and coherence among the extracted subgraphs.
(2) We adopt the text knowledge in these various fact verification datasets including FEVER (Thorne et al., 2018), Climate-Fever Diggelmann et al. (2020), Health-Fever (Sarrouti et al., 2021), COVID-FACT (Saakyan et al., 2021) and SCIFACT (Wadden et al., 2020a) to construct the data with vanilla pattern in our benchmark FactCHD. We only select the factual (supported) and non-factual (refuted) samples with their corresponding evidence and initially employ ChatGPT (GPT-3.5-turbo) to directly evaluate the claims in these datasets and select out samples where the model struggles to provide false answers.
Figure 3. Overview of the construction and pattern illustration of FactCHD.
3.2. Collect Realistic Data with Diverse Patterns
Prompt for Q&R Generation
We devise effective prompts to steer ChatGPT in generating the “Query-Response” contexts, comprising three vital elements: role description, factuality pattern, and factuality demonstration. The role description defines the system’s role and specifies the input and objective of the generation process. To ensure precise control over the type and quality of factuality-related samples, we give the explanation of corresponding factuality patterns and demonstrations to guide the model in generating “Query-Response” contexts.
3.3. Refine Prompts with Consistency
We utilize ChatGPT to generate “Query-Response” contexts, deploying manual prompts to enable its function as a role-playing agent, informed by specific demonstrative examples. Recognizing ChatGPT’s significant prompt sensitivity, we engage in an iterative refinement of them to ensure the generated data aligns closely with desired, realistic patterns, following the protocol delineated in § 3.1. Beginning with an initial fixed set of 100 examples, we select five samples for human similarity assessment, contrasting them with demonstrative examples. Employing the majority rule, we evaluate whether each context adheres to established similarity criteria. This iterative process allows for continuous prompt adjustment, assuring alignment between generated data patterns and desired aspirations.
3.4. Automated Screening for Enhanced Diversity
Aiming to improve the diversity of generated “Query-Response” contexts, we employ Sentence-BERT (SBERT)(Reimers and Gurevych, 2019) in an automated screening process, calculating average semantic similarity within contexts. This facilitates the identification and exclusion of highly similar samples, safeguarding dataset diversity. After the screening process, a total of 1,538 samples from the training set and 632 samples from the test set were removed, ensuring the final aggregate of samples is diverse. This meticulous elimination of semantically analogous entries ensures our collections exhibit varied and non-redundant queries and responses, enhancing the benchmark’s utility for evaluating diverse instances of fact-conflicting hallucination.
3.5. Generate Chain of Evidence
Our benchmark extends beyond merely identifying factual errors in “Query-Response” contexts produced by LLMs; it also necessitates the generation of coherent evidence chains to substantiate such judgments, all of which are rooted in factual knowledge and articulated by ChatGPT. Particularly, we exploit two types of knowledge sources (outlined in §3.1) as our factual foundation. ChatGPT employs either subgraph facts or textual facts to furnish comprehensive justifications when assigning appropriate labels to the “Query-Response” contexts. It’s pivotal to note that the efficacy of these explanations profoundly impacts the credibility and reliability of judgments, thereby fostering trust in predictive models. These explanatory outputs can also amplify users’ overall understanding.
3.6. Quality Filtering via Human Annotator
We devised filter rules encompassing three facets: pattern consistency, response factuality, and evidence chain logic, employing these to guide annotators through quality filtering. Our team of 21 annotators, each boasting proficient English reading skills and a minimum of undergraduate education, underwent uniform training per standardized annotation guidelines. The annotators leverage their own awareness and, when necessary, utilize search engines to ensure informed filtering decisions, discarding samples that flout established rules. Given the inherent subjectivity in defining matching quality via these rules, we divided both the annotators and samples into seven groups, each featuring three annotators reviewing the same data batch through a voting mechanism. Simultaneously judged mismatches by annotators led to sample discarding, resulting in final removal counts of 565 and 258 samples from the training and test sets, respectively.
4. Experiments
[…]

### v2

# FactCHD: Benchmarking Fact-Conflicting Hallucination Detection

@@
Abstract
Despite their impressive generative capabilities, LLMs are hindered by fact-conflicting hallucinations in real-world applications. The accurate identification of hallucinations in texts generated by LLMs, especially in complex inferential scenarios, is a relatively unexplored area. To address this gap, we present FactCHD, a dedicated benchmark designed for the detection of fact-conflicting hallucinations from LLMs. FactCHD features a diverse dataset that spans various factuality patterns, including vanilla, multi-hop, comparison, and set operation. A distinctive element of FactCHD is its integration of fact-based evidence chains, significantly enhancing the depth of evaluating the detectors’ explanations. Experiments on different LLMs expose the shortcomings of current approaches in detecting factual errors accurately. Furthermore, we introduce Truth-Triangulator that synthesizes reflective considerations by tool-enhanced ChatGPT and LoRA-tuning based on Llama2, aiming to yield more credible detection through the amalgamation of predictive results and evidence [^1] The benchmark is available at https://github.com/zjunlp/FactCHD..
1 Introduction
Large Language Models (LLMs) Zhao et al. (2023) are susceptible to generating text that, while seemingly credible, can be factually inaccurate or vague, leading to the spread of misinformation online Yin et al. (2023); Ji et al. (2023); Huang et al. (2023). This issue, referred to as fact-conflicting hallucination Zhang et al. (2023), arises from the incorporation of incorrect or obsolete knowledge into the models’ parameters and from the models’ inherent limitations in complex cognitive ability. These shortcomings constrain LLMs’ deployment in critical domains like finance, healthcare, and law, and amplify the propagation of erroneous information.
Figure 1: Illustration of fact-conflicting hallucination detection example from FactCHD, where the green part represents factual explanation core (body part) in the chain of evidence.
Therefore, it is crucial to effectively detect fact-conflicting hallucinations for mitigating or editing them Yao et al. (2023).
@@ 1 Introduction
However, traditional fact verification tasks Wadden et al. (2020a, b) are not suitable for LLM-based “Query-Response” data. Moreover, existing hallucination evaluation benchmarks Li et al. (2023); Muhlgay et al. (2023), predominantly centering on vanilla facts and textual content, lack in-depth exploration of complex operations among facts, thus rendering their coverage against fact-conflicting hallucinations suboptimal. To bridge this gap, we introduce an inherently rigorous, yet authentic, task scenario: fact-conflicting hallucination detection, devoid of explicit claims or evidence. As shown in Figure 1, when confronted with a query and its generated response, detectors, are impelled to harness both their intrinsic knowledge and external resources, while rendering a factual judgment accompanied by an elucidative explanation.
Datasets # Source Domains Construction Factuality Pattern Evaluation Expandability
Fever Thorne et al. (2018) Text General HG Van. Consist.
Climate-Fever Diggelmann et al. (2020) Text Climate Change HG Van. Consist.
Health-Fever Sarrouti et al. (2021) Text Health HG Van. Consist.
Sci-Fact Wadden et al. (2020a) Text Scientific HG Van. Consist.
CoVERT Mohr et al. (2022) Text COVID-19 HG Van. Consist.
TabFact Chen et al. (2020) Table General HG Van. Consist.
HoVer Jiang et al. (2020) Text General HG Mul. Consist.
FEVEROUS Aly and et al (2021) Text+Table General HG Van. Consist.
HaluEval Li et al. (2023) Text General MIX Van. Hallucination
FactCHD (ours) KGs (
) & Text
General&Vertical MIX Van.&Mul.&Com.&Set. Fact.+Chain.
Table 1: Comparison with existing fact-checking datasets. Our FactCHD include both general and vertical domains, such as health, COVID-19, climate, science, and medicine (genes, virus and disease). Van., Mul., Com. and Set. are abbreviations for the four distinct factuality patterns in FactCHD. We utilize the following abbreviations: HG - human-curated datasets; MIX - datasets containing both human-constructed and machine-generated data; Q&R - “Query-Response” context; Consist. - classification of consistency between claims and given evidence; Fact. - classification of factuality, with labels factual and non-factual; Chain. - chains of evidence.
In paving the way for future strides in hallucination evaluation, we introduce a new benchmark, Fact-Conflicting Hallucination Detection (FactCHD), tailored for LLMs and encompassing a variegated array of factuality patterns, including Vanilla, Multi-hops, Comparison, and Set-Operation patterns. For example in Figure 1, querying whether the “Alps” or “Pyrenees” are higher, exemplifies a comparison pattern, assessing the relative relationships among facts. Drawing inspiration from the adage “to know it and to know the reason why of it” by Zhuzi, FactCHD extends beyond mere “Query-Response” labeling of hallucinations, incorporating golden chains of evidence to assess if detectors can provide coherent explanations for factualness judgment. Acknowledging the challenges of collecting comprehensive data through exhaustive human annotation, we propose a scalable data construction approach that harnesses existing knowledge graphs (KGs) and textual knowledge to create simulated hallucination instances with ChatGPT, verified with human annotation, for the efficient development of FactCHD.
@@
We evaluate the performance of various LLMs (such as Alpaca, Llama2-chat, and ChatGPT) using our FactCHD benchmark across multiple settings: zero-shot, in-context learning, specialized detection tuning, and knowledge enhancement via retrieval/tools. The results indicate that specialized detection tuning and knowledge enhancement notably improve the detection of fact-conflicting hallucinations. Additionally, we present Truth-Triangulator framework grounded in “Triangulation” theory Valenza (2016). This system comprises three roles: ChatGPT, enhanced with tools as the Truth Seeker; a detect-specific expert based on Llama2-7B-LoRA as the Truth Guardian; and the Fact Verdict Manager, which amasses evidence from another role to fortify the reliability and accuracy of the derived conclusions. Truth-Triangulator emphasizes the use of cross-referencing generators to astutely evaluate and adjudicate responses with potential factual discrepancies. Key insights are summarized as:
  • We present FactCHD, a large-scale, multi-domain evaluation benchmark with diverse factual patterns and interpretable evidence chains, setting a new standard for detecting fact-conflicting hallucinations from LLMs.
  • We introduce a scalable data construction strategy leveraging KGs, etc. to efficiently develop a fact-conflicting hallucination dataset, offering stronger applicability due to the authentic, broad domain coverage of KGs.
  • We devise a triangulation-based framework Truth-Triangulator that employs cross-referencing generators for verifying LLM responses. Empirical results demonstrate the effectiveness of our approach in hallucination detection.
2 Related Work
@@
Figure 2: Overview of the factuality patterns invloved in our FactCHD.
Figure 3: Domain distribution of FactCHD, where “General-p” and “General-d” denote common domains derived from Wikipedia and Wikidata, respectively.
@@
Despite LLMs Zhao et al. (2023); Chen et al. (2022); OpenAI (2022) like ChatGPT demonstrating remarkable understanding and execution of user instructions, they are prone to confidently generating misleading “hallucinations” Ji et al. (2023); Ye et al. (2023); et al (2023a); Hu et al. (2023). These hallucinations, as categorized by Zhang et al. (2023), can be input-conflicting, context-conflicting, or fact-conflicting—with the latter being especially problematic due to the propagation of inaccurate factual information online. While previous studies have extensively examined hallucinations within natural language generation (NLG) for various NLP tasks Shuster et al. (2021); Creswell and Shanahan (2022); Das et al. (2023); Mallen et al. (2022), HaluEval Li et al. (2023) has emerged as a recent benchmark for assessing LLMs’ recognition of such errors. Different from HaLuEval which only analyzes ChatGPT’s ability to evaluate whether hallucinatory, we specifically focus on the evaluation of fact-conflicting hallucination by constructing an interpretable benchmark that can serve as a public platform for checking the factual errors of context with explanation.
Factuality Detection.
Our research is also related to prior works on fact verification in NLP tasks Thorne et al. (2018); Wadden et al. (2020b); Gupta et al. (2022); Dziri et al. (2022); Liang et al. (2023); Rashkin et al. (2021); Kryściński et al. (2020), expanding the understanding of factuality beyond binary judgments. The FRANK framework Pagnoni et al. (2021) offers a detailed typology of factual errors, while  Dhingra et al. (2019) explores lexical entailment in text generation. The FACTOR score by Muhlgay et al. (2023) evaluates LMs based on the likelihood of factual content. Existing benchmarks often miss the “Query-Response” context of LLMs, focusing solely on accuracy without providing explanatory rationales. Our contribution lies in: (1) presenting “Query-Response” formatted data for LLMs; (2) introducing metrics for interpretability in detecting fact-conflicting hallucinations; and (3) validating the effectiveness of Truth-Triangulator through cross-referential verification from multiple sources. The comparison with other datasets is detailed in Table 1.
3 Preliminaries
Task Formulation.
Detecting fact-conflicting hallucinations in LLMs involves discerning factual errors in responses to human queries. A comprehensive detector must not only classify responses as factual or non-factual but also provide explanations for its judgments. We define the task as follows: Input: A question $Q$ paired with an LLM-generated response $R$, which may contain various fact conflicts. Output: A combined label and explanation sequence $A=[l,e]$, where $l$ is the binary factuality label (factual or non-factual), and $e$ articulates the rationale behind the assigned label. We evaluate the quality of $e$ using the $ExpMatch$ metric through the golden evidence chains in FactCHD.
Factuality Patterns.
We aim to explore distinct patterns of factual errors in our FactCHD, vividly illustrated in Figure 2. These include the (1) vanilla pattern dealing with factual statements that can be objectively verified using established sources, the (2) multi-hops pattern involving the process of concluding by connecting multiple pieces of facts, the (3) comparison pattern referring to the act of evaluating and comparing relative worth and relations between different pieces of facts, and the (4) set-operation pattern involving manipulating and combining sets of elements using operations to analyze relations between different facts. We generated corresponding “Query-Response” examples based on these patterns.
Knowledge-Driven Factual Foundation for Reliable “Query-Response” Generation.
KGs Liang et al. (2024) serve as a rich trove of structured entities and relations, ideal for compositional reasoning and anchoring factual data. Alongside this, textual knowledge is critical for nuanced inference beyond basic facts. Our research focuses on collecting existing knowledge and integrating it into prompts as a factual foundation for “Query-Respons” and chain of evidence generation, as outlined below: (1) We use $438$ widespread relations from Wikidata Vrandecic and Krötzsch (2014) and PrimeKG Chandak et al. (2023) to create varied subgraphs via $K$-hop walks, forming the knowledge base for generating “Query-Respons” examples with multi-hop reasoning, fact comparison, and set operation patterns. (2) We utilize text knowledge from datasets such as FEVER Thorne et al. (2018), Climate-Fever Diggelmann et al. (2020), Health-Fever Sarrouti et al. (2021), COVID-FACT Saakyan et al. (2021), and SCIFACT Wadden et al. (2020a) for generating examples with vanilla pattern.
4 FactCHD Benchmark Construction
Building on the aforementioned preliminaries, we develop FactCHD, a dataset containing a wealth of training instances and an additional 6,960 carefully selected samples for evaluating fact-conflicting hallucinations from LLMs. Our dataset maintains a balanced representation of factual and non-factual categories, offering a robust framework for assessment. The statistics and domain distribution of FactCHD are depicted in Table 2 and Figure 3. Next, we outline the design principles of our benchmark as follows.
@@ 4 FactCHD Benchmark Construction
Table 2: Data Statistic of our FactCHD
4.1 Collect Realistic Data as Demonstration
Subsequently, adhering to the defined factuality patterns, we manually craft corresponding queries and utilize open-source LLMs (such as ChatGLM THUDM (2023), Alpaca, and Vicuna) to generate authentic responses. By manually annotating these responses for hallucinations, we acquire “Query-Response” examples annotated with hallucination presence. Our objective is to curate a robust dataset of hallucination cases that closely emulate real-world examples, providing a demonstrative foundation and establishing standards for prompt refinement to ensure alignment between generated hallucinations and these demonstrations.
4.2 Generate “Query-Response” Contexts
Figure 4: Overview of the construction process of FactCHD.
I want you to act as a data generator for a fallacy-finding task. You will be on the lookout for invalid arguments so you can call out any logical errors or inconsistencies that may be present in #Question# and #Answer#. I’ll tell you the right label: FACTUAL or NON-FACTUAL. Your job is to provide rational evidence-based feedback and point out any fallacies, faulty reasoning, false assumptions, or incorrect conclusions that may be present in the #Question# and #Answer#. Here are some examples:
#Golden Label#: NON-FACTUAL
#Query#: How many children did the leader of the country, where the distributor of "The Killer Inside Me" operates, have?
<Demonstrations>
#Golden Lable#: <Here is the right label for the factuality of the following sample > #Query#: <Here is the query for the LLM> #Response#: <Here is the response from the LLM according to the above query> #Evidence#: <Here is the evidence with fact to judge the factuality of above context>
#Output#:
Table 3: Instruction to generate evidence-based chains of facts for the “Query-Response” contexts. The role description is indicated by the yellow text, while the factuality demonstration is represented by the purple text.
Generate “Query-Response” based on Knowledge with ChatGPT.
We incorporate the knowledge collected on distinct factuality patterns into customized prompts, specifying whether to generate factual or non-factual responses. This guides ChatGPT in producing “Query-Respons” instances across various factuality categories with golden factuality labels. Detailed prompts are in Table 7 through  10 of  §A.
Refine Prompts with Consistency.
In the initial phase of generating “Query-Response” instances, we assess the consistency of five samples per pattern against demonstrations, using majority consensus. We subjectively evaluate each context’s adherence to the style and form of realistic demonstrations, employing iterative refinements of prompts to ensure that the generated data exhibits realistic patterns. We aim for an at least 95% consistency rate before scaling up the generation of “Query-Response” instances.
4.3 Automated Semantic Diversity Screening
To increase “Query-Response” context diversity, we employ Sentence-BERT (SBERT)Reimers and Gurevych (2019) to automatically compute semantic similarity matrices, filtering out near-duplicate samples to preserve dataset variety. During filtration, we removed 1,542 training and 832 test samples, guaranteeing a diverse final dataset. This careful pruning of semantically similar entries promotes a varied collection of queries and responses, bolstering the benchmark’s effectiveness for assessing diverse instances.
4.4 Generate Chain of Evidence
Our benchmark evaluates the detectors’ ability to not only identify hallucinations but also to provide effective explanations. It necessitates ChatGPT’s generation of coherent golden evidence chains grounded in factual knowledge for substantiating judgments. Utilizing subgraph or textual facts outlined in §section 3, along with the previously generated “Query-Response” pairs, ChatGPT delivers thorough justifications for labels assigned to “Query-Response” contexts. Detailed prompts are presented in Tables 3 These golden evidence chains are critical for assessing the explanatory validity of hallucination detectors.
4.5 Quality Filtering via Human Annotator
We craft filter rules to ensure pattern consistency, response factuality, and logical evidence chains for quality control in the annotation. 21 educated annotators are uniformly trained and utilize both their expertise and search tools for rigorous sample vetting. To minimize subjectivity, we divided them into seven triads, introducing a voting mechanism for data evaluation. Simultaneously judged mismatches by annotators led to sample discarding, resulting in final removal counts of 565 and 258 samples from the training and test sets, respectively. We involve Fleiss’s Kappa ($\kappa$) as a measure of inter-annotator agreement to assess the reliability of our annotations. We calculate $\kappa$ over a set of 100 annotated samples, resulting in a $\kappa$ value of 0.808, indicating substantial agreement.